OpenStudy (anonymous):

I need help tanx+secx=1

OpenStudy (anonymous):

I started with \[\tan x +\sec x -1=0\] \[\tan x+\frac{ 1 }{ \cos x }-1=0\] \[\frac{ \sin x }{ \cos x }+\frac{ 1 }{ \cos x }-1=0\] \[\frac{ \sin x+1 }{ \cos x }-1=0\] \[\frac{ \sin x+1 }{ \cos x }-\frac{ \cos x }{ \cos x }=0\] \[\frac{ \sin x+1-\cos x }{ \cos x }=0\] What do I do now or where did I go wrong?

OpenStudy (anonymous):

Original question was \[\tan x +\sec x =1\]